The sum of two numbers is 15 and their difference is 9. Find the two numbers.

Algebra ->  Coordinate Systems and Linear Equations  -> Lessons -> SOLUTION: Question 1: The sum of two numbers is 15 and their difference is 9. Find the two numbers.      Log On


   



Question 797126: Question 1: The sum of two numbers is 15 and their difference is 9. Find the two numbers.
Answer by mohanraj04(12)

given that
the sum of two numbers is 15 and difference of two numbers is 9
let X & Y be the two numbers
EQU:1 X+Y=15;
EQU:2 X-Y=9
ADD EQU:1 AND EQU:2
(X+Y)+(X-Y)=15+9
2X=24
X=24/2
X=12
substitute X=12 in equ:1
X+Y=15
12+Y=15
Y=15-12
Y=3
therefore the required two numbers are 12 & 3
to check
12+3=15
12-3=9